# Less is More: Improving LLM Alignment via Preference Data Selection

### v2

@@ 4.5 Ablation Study
A critical aspect of dataset filtering methods is their generalization capability - specifically, how well the method (selected subset) transfers to new models, similar optimization algorithms, etc. From this perspective, we examine DM’s robustness and effectiveness across several dimensions.
Data filtering remains effective for new architecture and datasets. While our initial experiments focused on the Llama architecture and tokenizers, we investigated DM’s effectiveness when applied to fundamentally different model architectures. We evaluate the Mistral-7B-Instruct model using its corresponding Mistral-UltraFeedback dataset, maintaining consistent experimental parameters with those described in Section 4.3. The AlpacaEval 2.0 evaluation results are presented in Table 4. Our experiments show that DM-MUL significantly outperforms baseline methods while maintaining output lengths (i.e. around 1600) comparable to the initial instruct model. We observe that models trained on the full dataset tend to generate notably longer responses (i.e., 4189), primarily due to their learned behavior of incorporating numerous hyperlinks (e.g., https://www.ibdb.com/) in an attempt to enhance response credibility. Similarly, models trained on the Low-Gap selected subset exhibit increased response lengths (i.e., 2525) and follow a comparable pattern of including hyperlinks. In contrast, models trained on DM-selected subsets do not display this behavior. We consider the inclusion of hyperlinks without proper RAG (Retrieval-Augmented Generation) (Lewis et al., 2020) or website verification mechanisms to be problematic, as such references can be unreliable.
@@
DM selected subsets are effective across other preference learning algorithms. We examine the effectiveness of our selected subsets across different DPO variants and compare the raw/LC win rates of models trained on full or selected subsets by IPO (Azar et al., 2024), KTO (Ethayarajh et al., 2024) and SLiC (Zhao et al., 2023). The results are presented in Figure LABEL:fig:abl_po. Our results demonstrate that DM-M consistently outperforms full-set training across all three algorithms, achieving particularly notable improvements in raw/LC win rates - over 12% - for the IPO algorithm. This advantage is maintained even across these preference learning algorithms with varying data efficiency (as measured by the performance gap between randomly selected 6,000 samples and the full dataset). These findings highlight the significant value of sample filtering for other preference learning. Additional results related to instruct model training can be found in Appendix D.4.
[…]
5 Related Work
Preference Learning Algorithms.Reinforcement Learning from Human Feedback has become a crucial component of recent Large Language Models (LLMs) such as ChatGPT (Ouyang et al., 2022). While the classical RLHF pipeline traditionally uses Proximal Policy Optimization, several alternative approaches have been proposed. These include but not limited other RL-based training algorithms (Li et al., 2023c; Zhong et al., 2024), rejection sampling (Dong et al., 2023; Gulcehre et al., 2023), conditional supervised fine-tuning (Lu et al., 2022; Yang et al., 2024; Zhang et al., 2024a), and Direct Preference Optimization (Rafailov et al., 2024). Among these alternatives, DPO has gained significant attention due to its simplicity and robust performance. Following the introduction of DPO, numerous works (Zhao et al., 2023; Azar et al., 2024; Ethayarajh et al., 2024; Meng et al., 2024; Tang et al., 2024; Han et al., 2024; Xu et al., 2024; Hong et al., 2024; Park et al., 2024) have attempted to improve its performance by modifying the DPO objective.
Data Selection in LLM Fine-tuning. Data selection is crucial in LLM post-training (Wang et al., 2024), primarily due to two key observations: LLM training typically converges rapidly, and excessive data can degrade model performance through overfitting or exposure to toxic content (Swayamdipta et al., 2020; Deng et al., 2023). Recent research has focused on enhancing instruction tuning efficiency by identifying high-quality subsets from large instruction datasets (Cao et al., 2024; Li et al., 2024; Xia et al., 2024), often adapting active learning query strategies (Ren et al., 2021) to assess sample uncertainty and diversity. However, data efficiency in preference learning remains relatively unexplored. While some studies studied annotation cost reduction in preference dataset creation (Muldrew et al., 2024; Yasunaga et al., 2024), our work firstly provides clear criteria for identifying informative samples while filtering toxic ones, thereby improving both DPO’s efficiency and performance. Furthermore, our method extends to iterative DPO (Xiong et al., 2024) and its variants (Zhang et al., 2024b), where training data is generated online.
@@ 6 Conclusion
Our research bridges the gap between algorithmic advances and data-focused approaches in Large Language Model (LLM) alignment by systematically examining how preference data quality affects Direct Preference Optimization (DPO). We address the issue of parameter shrinkage caused by noisy data and introduce a dual-margin-based strategy for selecting high-quality training examples. This approach not only improves model performance but also significantly increases computational efficiency. Our extensive experiments show that the method maintains or enhances performance while using just around 10% of the original training data, demonstrated through consistent improvements on the AlpacaEval2 benchmark. Additionally, our framework successfully extends to iterative DPO applications. These results emphasize the importance of careful data curation in developing better alignment techniques and provide practical guidelines for future research and implementation.
@@ Appendix A Datasets and Evaluation Details
Win rate judgments serve as a key evaluation metric across our experiments. For the TL;DR and HH, we slightly modify the evaluation prompts suggested by Rafailov et al. (2024) for the win rates judge. We use the original prompt designed for AlpacaEval and AlpacaEval 2.0 benchmark in all testing.
@@
FIRST provide a one-sentence comparison of the two summaries, explaining which you prefer and why. SECOND, on a new line, state only ”A” or ”B” to indicate your choice. Your response should use the format: Comparison: ¡one-sentence comparison and explanation¿ Preferred: ¡”A” or ”B”¿
@@
FIRST provide a one-sentence comparison of the two responses and explain which you feel is more helpful or harmless. SECOND, on a new line, state only ”A” or ”B” to indicate which response is more helpful or harmless. Your response should use the format: Comparison: ¡one-sentence comparison and explanation¿ More helpful or harmless: ¡”A” or ”B”¿
Appendix B More Implementation Details
@@
We present the implementation details of our baseline methods: TOP, MID, and BOT using margins calculated from three metrics, Conditional Perplexity (CPPL), External (Ex), and Implicit (Im) rewards. Subsequently, we describe the implementation of DM-MUL in our experimental setup.
@@
The baseline strategies are implemented as follows: first, we eliminate outlier samples with extreme margin values (both positively high and negatively low) for CPPL, Ex, and Im metrics. For the TOP and BOT strategies, we select samples based on their ranking positions at the upper and lower ends of the distribution, respectively. The MID strategy involves random sampling from the subset of samples whose margin values fall within the interval $[-\tau,\tau]$, where $\tau$ is set to 0.1 for CPPL and 1.0 for Ex/Im metrics.
@@
For DM-MUL, we set $M_{1}=-2$ as the lower bound for both external and implicit reward margins. The upper bound $M_{2}$ is determined dynamically based on two conditions: (1) The number of samples with margin values in the interval $[M_{2},\text{max margin}]$ is less than 30, or (2) The number of samples in $[M_{2},\text{max margin}]$ is less than max $\text{max margin}-M_{2}$.
Appendix C Visualization of Margin Distributions
C.1 Singular Margin Distribution
@@
The margin distributions calculated using CPPL margin, External and Implicit DPO reward margins, as illustrated in Figures LABEL:fig:cppl-margin,LABEL:fig:ex-margin,LABEL:fig:im-margin, reveal a notable concentration of sample margins around zero. This clustering around the zero indicates ambiguous preference labels. It leads to the challenge in preference learning, as evidenced by the substantially slower decrease in training loss (and slower increase in training margin) compared to samples with larger margins, as shown in Section 4.2.1.
C.2 Joint Margin Distribution
@@
To complement the left and middle subfigures in Figure LABEL:fig:visual, we present additional results showing the joint margin distributions of samples on the other datasets in Figure LABEL:fig:visual-extra. Our analysis reveals that external and implicit margins exhibit minimal correlation across all four datasets, while implicit margins calculated by different models maintain a high correlation. These further enhance the rationality of our design detail of DM: fusion of both margins and disentangling implicit margin from the target model (if the target model is a bit large and we want to accelerate the enumeration process of the full-set.)
Appendix D More Experimental Results
D.1 Train Loss and Margin Curves - 3B
@@
To complement the right subfigure in Figure LABEL:fig:visual, we present additional results showing the progression of training loss and margins throughout the DPO training process. The results are shown in Figure LABEL:fig:loss-margin. All strategies demonstrated consistent patterns in loss reduction: both TOP margin-oriented and BOT strategies achieved rapid decreases in training loss, while the MID strategy exhibited slower convergence and remained at significantly higher final loss values. Regarding training margins, TOP strategies achieved higher levels compared to BOT and MID approaches. Notably, our proposed DM-ADD and DM-MUL strategies demonstrated even larger margins than the Implicit Margin-TOP strategy.
D.2 Train Loss and Margin Curves - 8B
Figure LABEL:fig:loss-margin-8b presents the progression of training loss and margins during DPO training of the 8B model. We show the fine-tuned version of the Llama-3-8B (base) model trained on UltraFeedback as a representative case since all models exhibited similar trends, as in Figure LABEL:fig:loss-margin.
D.3 Resources and computation cost
@@
For all experiments, we utilized 8 A100 GPUs. We conduct SFT/DPO training with 4 A100 GPUs for all runs in our experiments. For both Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) training, we allocated 4 A100 GPUs per run. Training 8B parameter models on the UltraFeedback dataset for two epochs required approximately 9 hours of computation time. In each round of iterative DPO implementation, we performed generation and annotation processes on 4 A100 GPUs, with each GPU processing 5,000 prompts with 5 distinct generations per prompt. The overall generation that utilizes vLLM (Kwon et al., 2023) for acceleration takes about 1.5 hours, and the corresponding reward annotation takes about 2 hours.
D.4 More Results for Ablation Study on the DPO Variants
@@
As a complementary study to the results shown in Figure LABEL:fig:abl_po, we conducted experiments using the Llama-3-8B-Instruct model while maintaining all other experimental parameters. The results, presented in Figure LABEL:fig:abl_po_ins, demonstrate that models trained on subsets selected by DM-MUL achieved significantly higher win rates across most evaluation scenarios.
D.5 Ablation Study on Hyper-parameters
We investigate the impact of hyperparameters $\beta$ and training epochs on DPO performance. Specifically, we train the Llama-UltraFeedback model using DPO with $\beta=0.01$ for both one and two epochs. The model’s performance is evaluated using AlpacaEval 2.0 win rates, with results shown in Table 6. Our experimental results demonstrate that DM-MUL selected subsets significantly enhance the performance of vanilla DPO on these different hyperparameter configurations. Notably, DPO-DM-MUL achieves superior results compared to state-of-the-art preference learning algorithms such as SimPO (Meng et al., 2024), despite the latter utilizing the complete dataset for DPO training.
Table 6: Ablation study on $\beta$ ($\beta=0.01$): performance comparison of methods trained on Llama-Ultrafeedback dataset, benchmarked against state-of-the-art algorithms trained on the complete dataset.
Method LC (%) WR (%) Epochs
DPO 38.65 39.57 2
DPO 40.30 37.90 1
DPO-DM-MUL 43.66 42.52 2
DPO-DM-MUL 45.17 41.86 1
KTO 33.10 31.80 1
IPO 35.60 35.60 1
SimPO 43.72 39.88 1
@@ D.5 Ablation Study on Hyper-parameters
However, we observe that selecting a small $\beta$ value is not consistently a good choice for DPO training. Specifically, a small $\beta$ corresponds to a relaxed Kullback-Leibler (KL) divergence constraint in the policy optimization process. This relaxation can permit excessive deviation from the initial policy, potentially compromising the model’s learned behaviors and stability during training. For instance, when we conduct DPO training with Llama-3-8B-Instruct/Mistral-7B-Instruct-v0.2 on the Ex/Im-TOP selected 6,000 subsets from UltraFeedback, with $\beta=0.01$ and two epochs update, we find that although the model could respond normally to most questions, it sometimes outputs repeated or chaotic tokens, as shown in Figure 9. And their win rates on AlpacaEval 2.0 dropped by more than 10 points as a consequence.
@@
Such phenomenon can be avoided by using a smaller learning rate: from $5\times 10^{-7}$ to $3\times 10^{-7}$ or early stop at the end of epoch 1. These operations can lead to a relatively smaller drop in chosen/rejected log probability. However, we still do not recommend using small $\beta$, as more potential sacrifices in other tasks such as reasoning may also appear in this case.

### v3

@@ 4.4 Ablation Study
A critical aspect of dataset filtering methods is their generalization capability—specifically, the performance when transferred to new models or applied with similar optimization algorithms.
Data filtering remains effective for new LLM architecture.
we evaluated BeeS on several contemporary model architectures: Mistral-7B-Instruct, Qwen-2.5-7B-Instruct, and Qwen-2.5-14B-Instruct. Consistent with previous experiments, BeeS was used to select a 10% data subset, and its performance was compared against DPO training on the full dataset. The AlpacaEval 2.0 evaluation results are presented in the left panel of Figure 7. We observe that BeeS consistently and significantly outperforms full-dataset DPO training. Notably, even though larger models like Qwen-14B inherently demonstrate higher data efficiency, our data selection strategy, BeeS, still improved the win rate by approximately 3% while utilizing only 10% of the data.
@@
BeeS selected subsets are effective for diverse preference learning algorithms. We examine whether the subset selected by BeeS remains data efficient for DPO-variants like IPO (azar2024general, ), KTO (ethayarajh2024kto, ), and SLiC (zhao2023slic, ). We utilize the BeeS selected 6k-sample subset from Llama-UltraFeedback and the results are presented in the right panel of Figure 7. We observe that the high-margin subset consistently benefit these preference learning algorithms by outperforming full-set training. Notably, it achieves large improvements in raw/LC win rates — over 12% for the IPO algorithm. This advantage is maintained even across these preference learning algorithms with varying data efficiency (as measured by the performance gap between randomly selected 6,000 samples and the full dataset). These findings highlight the significant value of sample filtering for other preference learning. Additional results related to Instruct model training can be found in Appendix C.3.
Mistral-7B Qwen-2.5-7B Qwen-2.5-14B
LC (%) WR (%) LC (%) WR (%) LC (%) WR (%)
Initial 17.11 14.72 31.27 31.60 37.03 32.64
Full 18.00 18.77 39.67 38.24 49.99 46.81
BeeS 26.04 20.53 46.20 43.78 50.20 49.74
Figure 7: Ablation Study: (Left) different model choices (Mistral-7B-Instruct-v-0.2, Qwen-2.5-Instruct-7B and Qwen-2.5-Instruct-14B). BeeS selects a 6k-sample subset for training. (Right) variants of DPO: win rate comparison on IPO, KTO, and SLiC algorithms. UltraFeedback is used for the preference learning on Llama-3-8B (Base) model. Rand and BeeS select a 6k-sample subset.
@@ 5 Conclusion
Our research bridges the gap between algorithmic advances and data-focused approaches in Large Language Model (LLM) alignment by systematically examining how preference data quality affects Direct Preference Optimization (DPO). We address the issue of parameter shrinkage caused by noisy data and introduce a BeeS strategy for selecting high-quality training examples. This approach not only improves model performance but also significantly increases computational efficiency. Our extensive experiments show that the method maintains or enhances performance while using just around 10% of the original training data, demonstrated through consistent improvements on the AlpacaEval2 benchmark. Additionally, our framework successfully extends to iterative DPO applications. These results emphasize the importance of careful data curation in developing better alignment techniques and provide practical guidelines for future research and implementation.
@@ Appendix A Datasets and Evaluation Details
Win rate judgments serve as a key evaluation metric across our experiments. For the TL;DR and HH, we slightly modify the evaluation prompts suggested by [38] for the win rates judge. We use the original prompt designed for AlpacaEval and AlpacaEval 2.0 benchmark in all testing.
@@
FIRST provide a one-sentence comparison of the two summaries, explaining which you prefer and why. SECOND, on a new line, state only "A" or "B" to indicate your choice. Your response should use the format: Comparison: <one-sentence comparison and explanation> Preferred: <"A" or "B">
@@
FIRST provide a one-sentence comparison of the two responses and explain which you feel is more helpful or harmless. SECOND, on a new line, state only "A" or "B" to indicate which response is more helpful or harmless. Your response should use the format: Comparison: <one-sentence comparison and explanation> More helpful or harmless: <"A" or "B">
A.1 More Implementation Details
SFT. The SFT training of the Base model is carried out for two epochs with a learning rate of $2\times 10^{-5}$. Sample packing [45] is employed to accelerate the training, and we use a block size of 4096.
@@ A.1 More Implementation Details
We present the implementation details of our baseline methods: P, Z, and N using margins calculated from three metrics, IFD/Conditional Perplexity (CPPL), External (Ex), and Implicit (Im) rewards. Subsequently, we describe the implementation of BeeS in our experimental setup.
@@
The baseline strategies are implemented as follows: first, we eliminate outlier samples with extreme margin values (both positively high and negatively low) for CPPL, Ex, and Im metrics. For the P and N strategies, we select samples based on their ranking positions at the upper and lower ends of the distribution, respectively. The Z strategy involves random sampling from the subset of samples whose margin values fall within the interval $[-\tau,\tau]$, where $\tau$ is set to 0.1 for CPPL and 1.0 for Ex/Im metrics.
@@
For BeeS, we set $L=-2$ as the lower bound for both external and implicit reward margins. The upper bound $U$ is determined dynamically based on two conditions: (1) The number of samples with margin values in the interval $[U,\text{max margin}]$ is less than 30, or (2) The number of samples in $[U,\text{max margin}]$ is less than max $\text{max margin}-U$.
Appendix B Visualization of Margin Distributions
B.1 Singular Margin Distribution
@@
The margin distributions calculated using CPPL margin, External and Implicit DPO reward margins, as illustrated in Figures LABEL:fig:cppl-margin,LABEL:fig:ex-margin,LABEL:fig:im-margin, reveal a notable concentration of sample margins around zero. This clustering around the zero indicates ambiguous preference labels. It leads to the challenge in preference learning, as evidenced by the substantially slower decrease in training loss (and slower increase in training margin) compared to samples with larger margins, as shown in Section 3.2.
B.2 Joint Margin Distribution
@@
To complement the left and middle subfigures in Figure LABEL:fig:visual, we present additional results showing the joint margin distributions of samples on the other datasets in Figure LABEL:fig:visual-extra. Our analysis reveals that external and implicit margins exhibit minimal correlation across all four datasets, while implicit margins calculated by different models maintain a high correlation. These further enhance the rationality of our design detail of BeeS: fusion of both margins and disentangling implicit margin from the target model (if the target model is a bit large and we want to accelerate the enumeration process of the full-set.)
Appendix C More Experimental Results
C.1 Train Loss and Margin Curves - 3B
@@
To complement the right subfigure in Figure LABEL:fig:visual, we present additional results showing the progression of training loss and margins throughout the DPO training process. The results are shown in Figure LABEL:fig:loss-margin. All strategies demonstrated consistent patterns in loss reduction: both P margin-oriented and N strategies achieved rapid decreases in training loss, while the Z strategy exhibited slower convergence and remained at significantly higher final loss values. Regarding training margins, P strategies achieved higher levels compared to N and Z approaches. Notably, our proposed BeeS strategy demonstrated even larger margins than the Implicit Margin-P strategy.
C.2 Resources and computation cost
@@
For all experiments, we utilized 8 A100 GPUs. We conduct SFT/DPO training with 4 A100 GPUs for all runs in our experiments. For both Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) training, we allocated 4 A100 GPUs per run. Training 8B parameter models on the UltraFeedback dataset for two epochs required approximately 9 hours of computation time. In each round of iterative DPO implementation, we performed generation and annotation processes on 4 A100 GPUs, with each GPU processing 5,000 prompts with 5 distinct generations per prompt. The overall generation that utilizes vLLM [23] for acceleration takes about 1.5 hours, and the corresponding reward annotation takes about 2 hours.
C.3 More Results for Ablation Study on the DPO Variants
@@
As a complementary study to the results shown in Figure 7, we conducted experiments using the Llama-3-8B-Instruct model while maintaining all other experimental parameters. The results, presented in Figure LABEL:fig:abl_po_ins, demonstrate that models trained on subsets selected by BeeS achieved significantly higher win rates across most evaluation scenarios.
C.4 Hyperparameters Risks
@@
Although smaller $\beta$ can lead to higher data efficiency, we observe that it can bring potential issues for DPO training. Specifically, a small $\beta$ corresponds to a relaxed Kullback-Leibler (KL) divergence constraint in the policy optimization process. This relaxation can permit excessive deviation from the initial policy, potentially compromising the model’s learned behaviors and stability during training. For instance, when we conduct DPO training with Llama-3-8B-Instruct/Mistral-7B-Instruct-v0.2 on the Ex/Im-P selected 6,000 subsets from UltraFeedback, with $\beta=0.01$ and two epochs update, we find that although the model could respond normally to most questions, it sometimes outputs repeated or chaotic tokens, as shown in Figure 10. And their win rates on AlpacaEval 2.0 dropped by more than 10 points as a consequence.
@@
Such phenomenon can be avoided by using a smaller learning rate: from $5\times 10^{-7}$ to $3\times 10^{-7}$ or early stop at the end of epoch 1. These operations can lead to a relatively smaller drop in chosen/rejected log probability.
@@
Appendix D Limitations.
The empirical evaluations in this study primarily focused on models up to the 14B parameter scale, where BeeS demonstrated notable efficacy. Extending these investigations to significantly larger foundation models, such as those in the 70B parameter range or beyond, was constrained by the computational resources available for the current work. Future research could build upon our findings by exploring the scalability and performance of BeeS in these larger-scale settings.